OpenStudy (anonymous):

2. In the function f(x) = 4x how does an increase of 1 in the independent variable x affect f(x)?

OpenStudy (anonymous):

it increases f(x) by 4

OpenStudy (anonymous):

(x) will have an increase of 4 f(x) will have a decrease of 4 f(x) will have a decrease of 4x - 4 f(x) will have an increase of 4x + 4 are the choices

OpenStudy (anonymous):

reason being f(2) = (4)(2) or f(2+1) = 4(3)

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look at it f(x) = 4*x f(x+1) = 4(x+1) = 4x + 4

OpenStudy (anonymous):

thanks you very much

OpenStudy (anonymous):

you're welcome